# From Vulnerable to Resilient: Examining Parent and Teen Perceptions on How to Respond to Unwanted Cybergrooming Advances

### v1

@@ 1. Introduction
Today, 95% of U.S. teens have access to smartphones (Park, 2025) and over 93% use social media (DeAngelis, 2024), spending an average of 4.8 hours per day (DeAngelis, 2024), exposing them to risks such as cybergrooming (Whittle et al., 2014). These online risks have increased for minors; research has shown that over half of teens in the U.S. meet new people online (Lenhart, 2015), and 20% of minors have encountered online sexual solicitations from adults (Thorn, 2024a). Compared to more direct forms of online sexual exploitation, such as sextortion or sexual harassment, cybergrooming is a complex process of psychological manipulation that can extend over periods ranging from a few days to several years (De Santisteban and Gómez-Guadix, 2017). It involves a perpetrator who gradually befriended a teenager, strategically building trust and emotional bonds, and eventually sexually soliciting the teen either online or offline (Mladenović et al., 2021; Choo, 2009; Marchenko, 2017). Yet, research has also uncovered a dangerous “stranger danger” myth that frames unwanted online advances as originating from adult strangers (i.e., “predators”) who are imagined as overtly threatening or visibly “dangerous,” (Dedkova, 2015; Curran and Winther, 2025) when in reality, teens most often encounter unwanted sexual solicitations from seemingly innocuous individuals they meet online (Livingstone and Smith, 2014; Wolak et al., 2008) or already know in person (Alsoubai et al., 2022; Razi et al., 2020; Hartikainen et al., 2021). Regardless of who the perpetrator is, cybergrooming experiences have introduced severe consequences to minors. Children who have experienced online sexual solicitations may face severe physical consequences (e.g., unintended pregnancy, self-harm (NSPSS, 2025)) as well as mental health consequences (e.g., anxiety, depression, post-traumatic stress disorder (PTSD), and suicidal ideation (INHOPE, 2022; Schittenhelm et al., 2024)). They are also exposed to serious safety risks during offline meetups, including sexual assaults and even human trafficking (NSPSS, 2025; Front, 2023). These experiences can profoundly and lastingly affect teens’ overall well-being.
@@
A complementary alternative approach is to empower teens to recognize and manage online risks. While researchers have investigated sociotechnical approaches to support teens beyond surveillance (Gabrielli et al., 2020; Maeng and Lee, 2022; Rita and Shava, 2021; Piccolo et al., 2021), most chatbot-based interventions function as aftermath measures, serving as moderators to reduce the burden of in-person consulting. For example, Rita and Shava built an AI-based chatbot, named Report It, supporting teens searching resources or reporting cybergrooming incidents (Rita and Shava, 2021). We aim to explore ways to support sociotechnical educational interventions designed to empower teens to proactively recognize and cope with cybergrooming. This requires shifting attention away from perpetrator-centric views of detecting perpetrator behaviors toward teen-centric strategies to understand how parents and teens perceive vulnerability and protection.
This paper presents our first step toward that goal. We investigate how teens’ and parents’ thought processes work when responding to hypothetical yet realistic cybergrooming scenarios. By identifying recurring themes across their perceived vulnerable and protective responses, we provide insights into how parents and teens conceptualize online risk behaviors and risk-mitigating strategies. We study the perspectives of both teens and parents, as they may view and respond to the same online content differently. For example, teens may view random chats online as a friendly gesture, whereas parents believe that it is a threat to teens’ online safety (Wisniewski et al., 2017b). Yet, ensuring teens’ online safety requires collaborative efforts from both teens and parents, making it critical to understand how their perspectives align or diverge. Parents are often the primary stakeholders who care the most about their teens’ safety and are frequently responsible for making decisions about online risk management, monitoring, and protection (Akter et al., 2022, 2023a; Wisniewski et al., 2017a; Hashish et al., 2014; Ghosh et al., 2020). It is essential to understand if teens and parents perceive and respond to cybergrooming differently, and if so, how. These insights will contribute to designing effective and inclusive interventions. Thus, in this study, we ask the following research questions:
@@ 2.1. Understanding Cybergrooming Progression and Its Impacts on Teens
Understanding cybergrooming requires examining both how it is defined and how it unfolds behaviorally. Schittenhelm et al. defines cybergrooming as a process in which an adult uses digital communication tools to cultivate a deceptive relationship with a minor, ultimately aiming to facilitate sexual exploitation (Saha and De Choudhury, 2017). A literature review identified four recurring elements across definitions: perpetrators target minors, initiate contact through online technologies, pursue sexual objectives, and construct a pseudo-relationship to gain the teen’s trust (Schittenhelm et al., 2024). A widely used framework for understanding this progression is O’Connell’s six-stage taxonomy (O’Connell, 2003), developed through a five-year empirical study involving participant observation. The stages include: (1) Friendship Forming: perpetrators initiate contact and verify the teen’s identity through conversational exchanges, photos, or video calls; (2) Relationship Forming: perpetrators deepen familiarity by discussing personal topics such as family, school, and interests; (3) Risk Assessment: perpetrators evaluate whether the teen is isolated, ensure no external monitoring is occurring, and request secrecy; (4) Exclusivity: perpetrators cultivate emotional dependence by framing the relationship as unique or “special”; (5) Sexual: perpetrators gradually introduce sexualized topics, probe prior experiences, and normalize explicit conversations; (6) Conclusion: perpetrators transition to proposing offline meetings and coordinating logistics.
This taxonomy has shaped subsequent research on cybergrooming detection and analysis (Curran and Winther, 2025; Razi et al., 2023; Mladenović et al., 2021; Razi et al., 2021; Kloess et al., 2014; Bours and Kulsrud, 2019). For example, Gupta et al. (Gupta et al., 2012) identified the relationship-forming stage as most dominant in perpetrator conversations, while Razi et al. (Razi et al., 2023) used the taxonomy to classify online interactions as safe or unsafe. Gunawan et al. (Gunawan et al., 2016) further expanded this work by identifying 17 linguistic and behavioral features mapped to O’Connell’s stages, offering finer-grained insight into trust-building and escalation. Building on these stage-based and feature-based perspectives, our study examines how teens express both vulnerable and protective behaviors when encountering cybergrooming attempts.
@@ 3. Methodology
To address our research questions, we conducted an online survey study with 74 participants (51 parents and 23 teens) to examine how they responded to cybergrooming scenarios. In this section, we summarize the overview, scenario design, participant recruitment, ethical considerations, and analytic methods. Our study has been reviewed and approved by our Institutional Review Board (IRB).
@@ 3.2. Scenario Design
A key design consideration was ensuring that the scenarios felt realistic while avoiding unnecessary harm. Inspired by Agha et al. (Agha et al., 2024), we opted to use simulated cybergrooming scenarios to show the development of cybergrooming without exposing participants to real cybergrooming risks. We then asked participants to engage with the simulated cybergrooming scenarios. We framed the scenarios ambiguously by referring to the person in the scenarios as meeting “new people online,” rather than other explicit terms (e.g., “perpetrator”, “predator”), highlighting that the goal is to engage safely while preventing unwanted online sexual solicitation. Therefore, we did not frame each scenario as a definitive cybergrooming attempt. As such, the online scenario involving a stranger, i.e., a person they do not know, was used to debunk the stranger-danger myth (Curran and Winther, 2025; Dedkova, 2015) and to place them in a situation where a stranger can seem friendly and not appear unsafe. These scenarios also reflect contemporary patterns of online socialization among youth, in which interacting with new people online is far more common than many adults expect.
@@
S1 Asking for phone number Friendship Forming The online stranger is asking for the teen’s phone number
S2 Asking for family information Relationship Forming The online stranger is requesting the teen’s family information and relationship
S3 Acknowledging wrongdoing Risk Assessment The online stranger is acknowledging to the teen that chatting with an underage is wrong
S4 Asking if alone Risk Assessment The online stranger is asking if the teen will be alone on the weekend
S5 Asking about romantic relationship Exclusivity The online stranger is asking about the teen’s current romantic relationship
S6 Building mutual trust Exclusivity The online stranger is trying to show that they are special to the teen
S7 Asking for sensitive pictures Sexual The online stranger is complimenting the teen’s body and asking for their hot pics
S8 Asking about sexual experiences Sexual The online stranger is asking about the teen’s previous sexual experiences and their willingness to have sex
S9 Ice-cream fight Sexual The online stranger is flirting with the teen
S10 Offline meetup request Conclusion The online stranger is offering an offline meetup to the teen
@@ 3.5. Data Analysis
For the quantitative analysis, we calculated corresponding percentages at both the theme- and subtheme-levels (see Tables 5, 6). We then applied chi-square ($\chi^{2}$) analyses with adjusted standardized residuals in an exploratory manner to examine distributional variations in participants’ perceived vulnerable behaviors and protective strategies across cybergrooming scenarios (see Figure 3, 4).
@@ 4.1. Descriptive Statistics of the Participants
We recruited 74 participants in total: 51 parents and 23 teens. They represent diverse backgrounds in terms of age, gender, sexual orientation, race, and daily social media use. Additionally, parent participants reported their highest level of education, household income, and marital status. See Table 8 for details. Drawing from pre-validated survey questions from the Youth Internet Safety Survey (YISS) (Al-Salehi et al., 2021), Figure 2 provides an overview of parents’ perceptions and teens’ self-reports regarding prior exposure to unwanted sexual solicitations on social media. This figure demonstrates whether teens (self-reported or reported by their parents) had previously encountered such solicitations and how frequently these experiences occurred. Across all three types of solicitations, teens consistently reported higher levels of prior exposure than parents. Most teens who had encountered such solicitations reported that these experiences occurred once to a few times per year, with a notable number reporting more frequent exposure. In contrast, parents’ reports were more concentrated in the “not at all” or “I don’t know” categories, indicating that parents may underestimate the frequency with which their teens experience such solicitations.
@@ 5.4. Limitation and Future Work
While our study provided some insights into how parents and teens responded to cybergrooming, several limitations should be acknowledged and addressed in future research. First, our participant pool included only parents and teens, excluding other stakeholders such as child safety experts. Future research should broaden recruitment to include experts to capture a more holistic perspective. Second, although we grounded our scenarios in authentic online stranger and teen conversations from the Perverted-Justice (PJ) dataset, they were still hypothetical exercises. Future studies could incorporate anonymized in situ conversational data to capture real-world interactions better. Third, while we identified a stage-based taxonomy of protective strategies, we did not evaluate their effectiveness in combating cybergrooming or preventing escalation. Future research could use controlled experiments—such as simulations with large language models (LLMs)—to test how online strangers will respond to such protective responses and which strategies most effectively push back grooming advances. Fourth, due to the sample size, we are unable to do a comparison between teens of different ages. Future work could collect a large sample size with teens of different ages and examine the differences. Lastly, future work could explore opportunities to strengthen post-incident responses and detection systems. For instance, detection algorithms with supporting teams and agencies, rather than existing methods, may offer safer and more effective pathways for intervention. This direction extends beyond the scope of this study but represents an important area for continued exploration. Last, our vignette-based scenarios primed participants regarding the potential for unwanted sexual solicitations and the desire to prevent them to orient them to the task. As a result, our findings reflect how parents and teens reason about vulnerability and protection after attuned to the risk, rather than how teens gradually detect cybergrooming cues in real time. Therefore, there is room for future work to study how teens gradually detect and interpret cybergrooming cues in real time, using longitudinal, in situ, ecologically valid, and developmentall appropriate methods.
@@ Acknowledgements.
This research is supported by the U.S. National Science Foundation under grants #CNS-2330940, #CNS-2550834, #IIS-2550812, #TI-2550746, and by the William T. Grant Foundation grant #187941. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the research sponsors.

### v2

# arXiv:2601.21518v2[cs.HC] 28 Feb 2026

@@ 1. Introduction
Today, 95% of U.S. teens have access to smartphones (Park, 2025) and over 93% use social media (DeAngelis, 2024), spending an average of 4.8 hours per day (DeAngelis, 2024), exposing them to risks such as cybergrooming (Whittle et al., 2014). These online risks have increased for minors; research has shown that over half of teens in the U.S. meet new people online (Lenhart, 2015), and 20% of minors have encountered online sexual solicitations from adults (Thorn, 2024a). Compared to more direct forms of online sexual exploitation, such as sextortion or sexual harassment, cybergrooming is a complex process of psychological manipulation that can extend over periods ranging from a few days to several years (De Santisteban and Gómez-Guadix, 2017). It involves a perpetrator who gradually befriended a teenager, strategically building trust and emotional bonds, and eventually sexually soliciting the teen either online or offline (Mladenović et al., 2021; Choo, 2009; Marchenko, 2017). Yet, research has also uncovered a dangerous “stranger danger” myth that frames unwanted online advances as originating from adult strangers (i.e., “predators”) who are imagined as overtly suspicious or visibly “dangerous,” (Dedkova, 2015; Curran and Winther, 2025) when in reality, teens most often encounter unwanted sexual solicitations from seemingly innocuous individuals they meet online (Livingstone and Smith, 2014; Wolak et al., 2008) or already know in person (Alsoubai et al., 2022; Razi et al., 2020; Hartikainen et al., 2021). Regardless of who the perpetrator is, cybergrooming experiences have introduced severe consequences to minors. Children who have experienced online sexual solicitations may face severe physical consequences (e.g., unintended pregnancy, self-harm (NSPSS, 2025)) as well as mental health consequences (e.g., anxiety, depression, post-traumatic stress disorder (PTSD), and suicidal ideation (INHOPE, 2022; Schittenhelm et al., 2024)). They are also exposed to serious safety risks during offline meetups, including sexual assaults and even human trafficking (NSPSS, 2025; Front, 2023). These experiences can profoundly and lastingly affect teens’ overall well-being.
@@
A complementary alternative approach is to empower teens to recognize and manage online risks. While researchers have investigated sociotechnical approaches to support teens beyond surveillance (Gabrielli et al., 2020; Maeng and Lee, 2022; Rita and Shava, 2021; Piccolo et al., 2021), most chatbot-based interventions function as aftermath measures, serving as moderators to reduce the burden of in-person consulting. For example, Rita and Shava built an AI-based chatbot, named Report It, supporting teens searching resources or reporting cybergrooming incidents (Rita and Shava, 2021). We aim to explore ways to support sociotechnical educational interventions designed to empower teens to proactively recognize and cope with cybergrooming. This goal requires shifting attention away from perpetrator-centric views of detecting perpetrator behaviors toward teen-centric strategies to understand how parents and teens perceive vulnerability and protection.
This paper presents our first step toward that goal. We investigate how teens’ and parents’ thought processes work when responding to hypothetical yet realistic cybergrooming scenarios. By identifying recurring themes across their perceived vulnerable and protective responses, we provide insights into how parents and teens conceptualize online risk behaviors and risk-mitigating strategies. We study the perspectives of both teens and parents, as they may view and respond to the same online content differently. For example, teens may view random chats online as a friendly gesture, whereas parents believe that it is a threat to teens’ online safety (Wisniewski et al., 2017b). Yet, ensuring teens’ online safety requires collaborative efforts from both teens and parents, making it critical to understand how their perspectives align or diverge. Parents are often the primary stakeholders who care the most about their teens’ safety and are frequently responsible for making decisions about online risk management, monitoring, and protection (Akter et al., 2022, 2023a; Wisniewski et al., 2017a; Hashish et al., 2014; Ghosh et al., 2020). It is essential to understand if teens and parents perceive and respond to cybergrooming differently, and if so, how. These insights will contribute to the design of effective and inclusive interventions. Thus, in this study, we ask the following research questions:
@@ 2.1. Understanding Cybergrooming Progression and Its Impacts on Teens
Understanding cybergrooming requires examining both how it is defined and how it unfolds behaviorally. Schittenhelm et al. defines cybergrooming as a process in which an adult uses digital communication tools to cultivate a deceptive relationship with a minor, ultimately aiming to facilitate sexual exploitation (Saha and De Choudhury, 2017). A literature review identified four recurring elements across definitions: perpetrators target minors, initiate contact through online technologies, pursue sexual objectives, and construct a pseudo-relationship to gain the teen’s trust (Schittenhelm et al., 2024). A widely used framework for understanding this progression is O’Connell’s six-stage taxonomy (O’Connell, 2003), developed through a five-year empirical study involving participant observation. The stages include: (1) Friendship Forming: perpetrators initiate contact and verify the teen’s identity through conversational exchanges, photos, or video calls; (2) Relationship Forming: perpetrators deepen familiarity by discussing personal topics such as family, school, and interests; (3) Risk Assessment: perpetrators evaluate whether the teen is isolated, ensure no external monitoring is occurring, and request secrecy; (4) Exclusivity: perpetrators cultivate emotional dependence by framing the relationship as unique or “special”; (5) Sexual: perpetrators gradually introduce sexualized topics, probe prior experiences, and normalize explicit conversations; (6) Conclusion: perpetrators transition to proposing offline meetings and coordinating logistics. This taxonomy has shaped subsequent research on cybergrooming detection and analysis (Curran and Winther, 2025; Razi et al., 2023; Mladenović et al., 2021; Razi et al., 2021; Kloess et al., 2014; Bours and Kulsrud, 2019). For example, Gupta et al. (Gupta et al., 2012) identified the relationship-forming stage as most dominant in perpetrator conversations, while Razi et al. (Razi et al., 2023) used the taxonomy to classify online interactions as safe or unsafe. Gunawan et al. (Gunawan et al., 2016) further expanded this work by identifying 17 linguistic and behavioral features mapped to O’Connell’s stages, offering finer-grained insight into trust-building and escalation. Building on these stage-based and feature-based perspectives, our study examines how teens express both vulnerable and protective behaviors when encountering cybergrooming attempts.
@@ 3. Methodology
To address our research questions, we conducted an online survey study with 74 participants (51 parents and 23 teens) to examine how they responded to cybergrooming scenarios. In this section, we present the overview, scenario design, participant recruitment, ethical considerations, and analytic methods. Our study has been reviewed and approved by our Institutional Review Board (IRB).
@@ 3.2. Scenario Design
A key design consideration was ensuring that the scenarios felt realistic while avoiding unnecessary harm. Inspired by Agha et al. (Agha et al., 2024), we opted to use simulated cybergrooming scenarios to show the development of cybergrooming without exposing participants to real cybergrooming risks. We asked participants to engage with the simulated cybergrooming scenarios. We framed the scenarios ambiguously by referring to the person in the scenarios as meeting “new people online,” rather than in explicit terms (e.g., “perpetrator,” “predator”), highlighting that the goal is to engage safely while preventing unwanted online sexual solicitation. Therefore, we did not frame scenarios as definitive cases of cybergrooming. As such, the scenario involving an online stranger, i.e., a person they do not know, was used to debunk the stranger-danger myth (Curran and Winther, 2025; Dedkova, 2015) and to place them in a situation where a stranger can seem friendly and not appear unsafe. These scenarios also reflect contemporary patterns of youth online socialization, where interacting with new people online is more common than many adults expect.
@@
S1 Asking for phone number Friendship Forming The online stranger is asking for the teen’s phone number.
S2 Asking for family information Relationship Forming The online stranger is requesting the teen’s family information and relationship.
S3 Acknowledging wrongdoing Risk Assessment The online stranger is acknowledging to the teen that chatting with an underage is wrong.
S4 Asking if alone Risk Assessment The online stranger is asking if the teen will be alone on the weekend.
S5 Asking about romantic relationship Exclusivity The online stranger is asking about the teen’s current romantic relationship.
S6 Building mutual trust Exclusivity The online stranger is trying to show that they are special to the teen.
S7 Asking for sensitive pictures Sexual The online stranger is complimenting the teen’s body and asking for their hot pics.
S8 Asking about sexual experiences Sexual The online stranger is asking about the teen’s previous sexual experiences and their willingness to have sex.
S9 Ice-cream fight Sexual The online stranger is flirting with the teen.
S10 Offline meetup request Conclusion The online stranger is offering an offline meetup to the teen.
@@ 3.5. Data Analysis
For the quantitative analysis, we calculated corresponding percentages at both the theme- and subtheme-levels (see Tables 5, 6). We applied chi-square ($\chi^{2}$) analyses with adjusted standardized residuals in an exploratory manner to descriptively examine distributional variations in participants’ perceived vulnerable behaviors and protective strategies across cybergrooming scenarios. Because responses were collected from the same participants across multiple scenarios, these analyses are intended for pattern exploration rather than confirmatory inference. Adjusted standardized residuals were reported as descriptive diagnostics.
@@ 4.1. Descriptive Statistics of the Participants
We recruited 74 participants in total: 51 parents and 23 teens. They represent diverse backgrounds in terms of age, gender, sexual orientation, race, and daily social media use. Additionally, parent participants reported their highest level of education, household income, and marital status. See Table 8 for details. Drawing from pre-validated survey questions from the Youth Internet Safety Survey (YISS) (Al-Salehi et al., 2021), Figure 2 provides an overview of parents’ perceptions and teens’ self-reports regarding prior exposure to unwanted sexual solicitations on social media. This figure demonstrates that more than half of the teens (based on self-reported values) had previously encountered such solicitations. Across all three types of solicitations, teens consistently reported higher levels of prior exposure than parents. Most teens who had encountered such solicitations reported that these experiences occurred once to a few times per year, with a notable number reporting more frequent exposure. In contrast, parents’ reports were more concentrated in the “not at all” or “I don’t know” categories, indicating that parents may underestimate the frequency with which their teens experience such solicitations.
@@ 5.4. Limitation and Future Work
While our study provided some insights into how parents and teens responded to cybergrooming, several limitations should be acknowledged and addressed in future research. First, our participant pool included only parents and teens, excluding other stakeholders such as child safety experts. Future research should broaden recruitment to include experts to capture a more holistic perspective. Second, although we grounded our scenarios in authentic online stranger and teen conversations from the Perverted-Justice (PJ) dataset, they were still hypothetical exercises. Future studies could incorporate anonymized in situ conversational data to capture real-world interactions better. Third, while we identified a stage-based taxonomy of protective strategies, we did not evaluate their effectiveness in combating cybergrooming or preventing escalation. Future research could use controlled experiments—such as simulations with large language models (LLMs)—to test how online strangers will respond to such protective responses and which strategies most effectively push back grooming advances. Fourth, due to the sample size, we are unable to do a comparison between teens of different ages. Future work could collect a large sample size with teens of different ages and examine the differences. Lastly, future work could explore opportunities to strengthen post-incident responses and detection systems. For instance, detection algorithms with supporting teams and agencies, rather than existing methods, may offer safer and more effective pathways for intervention. This direction extends beyond the scope of this study but represents an important area for continued exploration. Last, our vignette-based scenarios primed participants regarding the potential for unwanted sexual solicitations and the desire to prevent them, to orient them to the task. As a result, our findings reflect how parents and teens reason about vulnerability and protection after being attuned to the risk, rather than how teens gradually detect cybergrooming cues in real time. Therefore, there is room for future work to study how teens gradually detect and interpret cybergrooming cues in real time, using longitudinal, in situ, ecologically valid, and developmentally appropriate methods.
@@ Acknowledgements.
This research is supported by the U.S. National Science Foundation under grants #CNS-2330940, #CNS-2550834, #IIS-2550812, #TI-2550746, by the William T. Grant Foundation grant #187941, and by the Center for Human-Computer Interaction (CHCI) at Virginia Tech. The views expressed in this material are those of the authors and do not necessarily reflect the views of the funding agencies.